# Study of transport properties of a hot and dense QCD matter using a novel approximation method

### v1

@@ Abstract
We have studied the charge and heat transport properties of a hot and dense QCD matter using a novel approximation method within the quasiparticle model. Utilizing a novel collision integral for both the relaxation time approximation (RTA) and the Bhatnagar-Gross-Krook (BGK) models, we have solved the relativistic Boltzmann transport equation to estimate the electrical conductivity and the thermal conductivity. We have investigated the temperature dependence of these transport coefficients. We have also provided a comparison between our findings and those of the standard RTA and BGK models. Further, we have explored the temperature dependence of the thermal diffusion constant and the Lorenz number using the novel approaches of the aforesaid models.
@@ 1 Introduction
A long time goal of the physics of ultrarelativistic heavy ion collisions is the observation of new states of matter predicted by the quantum chromodynamics (QCD). The experimental evidences of the relativistic heavy ion collider (RHIC) [1] at BNL and the large hadron collider (LHC) [2] at CERN indicate a new state of strongly interacting matter known as quark-gluon plasma (QGP). Following the discovery of QGP, various properties of the strongly coupled QCD matter became a topic of high interest. Different research groups have studied different features of QGP in recent years, such as the thermodynamic properties [3, 4], the relativistic hydrodynamics [5, 6, 7], the heavy quark diffusion in QCD matter [8, 9, 10], the memory effect on heavy quark dynamics [11, 12], the dissipative spin dynamics in relativistic matter [13, 14] etc. Nowadays, there is much interest in the phase diagram region at finite temperature and baryon chemical potential [15, 16]. Many observations have also been made recently to investigate the impact of a strong magnetic field on the characteristics of hot QCD matter, such as the chiral magnetic effect [17, 18], the chiral vortical effect [19], the axial Hall current [20], the refractive indices and decay constant [21, 22], the photon and dilepton productions from QGP [23, 24, 25], the thermodynamic and magnetic properties [26, 27, 28, 29, 30, 31], the axial magnetic effect [32, 33], the thermoelectric responses of hot QCD medium in a time varying magnetic field [34, 35], the magnetic and inverse magnetic catalysis [36, 37], the nonlinear electromagnetic current [38, 39] and the magnetohydrodynamics [40, 41].
@@
It is crucial to investigate different transport coefficients since they shed light on the genesis of the hot QCD medium. Among these transport coefficients, the thermal conductivity () and the electrical conductivity () characterize the movement of heat and charges in the medium, respectively. The hydrodynamic development of the strongly interacting matter at finite baryon density also depends on these two conductivities [42, 43]. Therefore, significant effects on the observables at heavy ion collisions might result from changes in the charge and heat transport coefficients. There are several methods for determining these conductivities, such as the relativistic Boltzmann transport equation [44, 45, 46], the Chapman-Enskog approximation [47, 48], the lattice simulation [49, 50, 51], the correlator technique using Green-Kubo formula [52, 53, 54] etc. Previous research has examined the impact of magnetic field on the electrical conductivity using a variety of methods, including the diluted instanton-liquid model [52], the quenched SU(2) lattice gauge theory [55], the effective fugacity technique [56], the diagrammatic method utilizing the real-time formalism [57] etc. Studies on the electrical and thermal conductivities of hot QCD matter in the presence of magnetic field within the kinetic theory approach have been carried out in references [58, 59, 60, 61].
@@
The application of theoretical methodologies intended for systems of this type has been made viable by the strongly coupled fluid behavior of QGP. A simplified version of the kinetic equation with a simplified form of the collision term is frequently used because of various technical limitations. It is referred as the relaxation time approximation (RTA) [62]. In this model, the mean-free approach reduced the complexity of the collision term. The development in the nonrelativistic case has already been achieved by Bhatnagar-Gross-Krook (BGK) model [63] and Welander model [64]. This approximation is utilized in many domains of physics and is recently used to investigate the hydrodynamization of the matter generated by the ultrarelativistic collisions of heavy ions. The relaxation time approximation has become essential for exploring the domain and microscopic basis of the relevance of hydrodynamics with relativistic effects. Although widely used, the approximation presented in ref. [62] has some fundamental challenges. It conflicts with some macroscopic and microscopic conservation laws. Recently, a paper [65] provided a novel relaxation time approximation that fixes such fundamental issues and maintains the core features of the linearized Boltzmann collision operator. In light of the momentum-dependent relaxation time, the novel relaxation time approximation is significant as it maintains the fundamental collision features associated with the microscopic conservation laws. They showed how this adjustment impacts the particle diffusion and the bulk viscosity. In a recent study [66], the transport coefficients related to charge and heat for a rotating QGP medium have been derived within the RTA model utilizing the aforementioned collision integral.
@@
In this study, we estimate the response functions of charge flow and heat flow, such as the electrical conductivity () and the thermal conductivity (), to investigate the charge and the thermal transport properties of the hot and dense QCD matter. These transport coefficients have been computed using a modified collision integral in the relativistic Boltzmann transport equation within the kinetic theory approach. We also explore the thermal diffusion constant and the Lorenz number as the applications of charge and heat transport properties in an analogous environment. The thermal diffusion constant shows the overall rate of heat transport in the medium, while the Lorenz number illustrates the connection between the thermal and electrical conductivities. We use the quasiparticle description of partons, where the rest masses are replaced by the masses generated in the medium. The quark-gluon plasma is viewed as a medium made up of thermally massive quasiparticles in this framework [67, 68].
@@
The rest of the paper has been organized as follows: Using the modified collision integral, the charge and heat transport coefficients from the relativistic Boltzmann transport equation are calculated in section 2. In section 3, we have discussed the results of the aforesaid transport coefficients using the quasiparticle description of partons. Section 4 describes various applications and their results obtained from these coefficients, such as the Lorenz number and the thermal diffusion constant. In section 5, the results are summarized.
@@ 2 Charge and heat transport properties of a hot and dense QCD medium
where F is the force field acting on the particles in the medium and is the mass of the particle. The term on the right-hand side results from particle collisions within the medium. This term’s zero value denotes the Vlasov equation, which describes a collisionless system. The intricate character of the collision term presents a number of challenges even for the simplest solution of the above eq. (1). The term represents the instantaneous change in the distribution function due to collisions between particles. In order to arrive at a comprehensive solution, we will first go over a few mathematical models that allow us to address the collision term in a suitable manner.
@@
where we can write and is the relaxation time for collisions, i.e. the time required to bring the perturbed system back to its equilibrium state. We can suppose the quark distribution function is near equilibrium with a slight divergence from equilibrium in order to permit linearization of the Boltzmann equation. Here, the term acts as a damping frequency. The initial issue with this model seems to be the immediate conservation of charge. To remove the crunch, the Bhatnagar-Gross-Krook (BGK) model presents a new type of collision kernel, denoted as
where represents the equilibrium density and is known as the perturbed density or the fluctuating density. The perturbed density is calculated as which gives zero value after integrating over momentum, thus indicating instantaneous conservation of the particle number during collisions in the system. Such fundamental problems are resolved and the essential characteristics of the linearized Boltzmann collision operator are preserved via a novel relaxation time approximation that was recently presented in a research [65].
Here, we are considering the general relativistic covariant form of the relativistic Boltzmann transport equation for an isotropic medium of quarks, antiquarks and gluons, which is given as
where is the electromagnetic field strength tensor. In order to see the effects of electric field, we take only and and vice versa, components of the electromagnetic field strength tensor, i.e. and in our calculation, thus the RBTE (4) takes the following form,
@@ 2.1 Electrical conductivity in the novel RTA and the novel BGK models
The electric current density of the medium must be determined in order to see how the electric field affects the charge transport phenomena. An external electric field can cause an infinitesimal disturbance to a thermal QCD medium that contains quarks, antiquarks, and gluons of distinct flavors. Due to this electric field, there is an induced current density in the medium, which is given by
@@
By comparing eq. (6) and eq. (7), one can obtain the expression of the electrical conductivity.
By solving the relativistic Boltzmann transport equation, the infinitesimal changes in the quark and antiquark distribution functions can be obtained. The relativistic Boltzmann transport equation in the novel-RTA model can be written as
@@
respectively, where , , and represents the chemical potential of flavor of quark. Getting inspired by [69, 70, 71], we have used the following expression of the relaxation time,
where is an arbitrary constant which controls the energy dependence of the relaxation time and the expression for the is given [72] by
Here, is the QCD running coupling constant, which is a function of both temperature and chemical potential and it has the following [73] form,
where GeV, for electrically charged particles (quarks and antiquarks) and for gluons. Using the expression of in eq. (8), we have
@@
In the novel-BGK model, the relativistic Boltzmann transport equation can be expressed as
where
The infinitesimal change in the quark distribution function is obtained (details are given in appendix B) as
@@ 2.2 Thermal conductivity in the novel RTA and the novel BGK models
Heat flow results from the deviation of the system from its equilibrium condition caused by the presence of a temperature gradient. Through a proportionality constant, i.e. the thermal conductivity, it is directly proportional to the temperature gradient. The calculation of the thermal conductivity is necessary in examining the heat transmission in a medium. In four-vector notation, the heat flow is written as
@@
where , and represent the energy density, the pressure and the particle number density, respectively, which are respectively given by the following equations,
@@
Now, solving this equation one can extract the expression of (for a detailed calculation see appendix C) as
Similarly for antiquarks,
@@
In the novel-BGK model, the relativistic Boltzmann transport equation takes the following form,
@@
The infinitesimal change in the quark distribution function is obtained (for a detailed calculation see appendix D) as
Similarly, is obtained as
@@ 3 Results and discussions
respectively, with , where the expression of is given in eq. (18). It can be seen from eq. (18) that, depends on both temperature and chemical potential. In this work, we have used GeV. The accessible energy scales for a thermal QCD medium are related to the temperature, quark mass and chemical potential. The QGP phase, which is made up of light quarks, light antiquarks and gluons, has a temperature high enough to reach the maximum energy scale. The hard thermal loop (HTL) perturbation theory is one of the effective theories that is applied to tackle the divergences that arise in the computation of the QCD thermodynamic observables, the transport coefficients and the amplitudes in the high temperature regime.
In this section, the charge and heat transport coefficients are investigated with the thermal masses as given in equations (55) and (56). In the novel RTA and the novel BGK models, quantities and appear in the denominators of the electrical conductivity and the thermal conductivity expressions. Thus the knowledge on the variations of these quantities with temperature and momentum is important to understand the variations of the conductivities. In figures 1 and 2, we have plotted and for momentum-dependent relaxation time () as well as for momentum-independent relaxation time (), respectively. We have found decreasing trends of and with the increase of temperature (figures 1(a) and 2(a)) and their increasing trends with the increase of momentum (figures 1(b) and 2(b)). In both the cases, it is observed that the momentum dependence of the relaxation time enhances both and . Thus, through the effects of the momentum-dependent relaxation time on and , it is possible to comprehend how the momentum-dependent relaxation time would modulate the transport coefficients.
Figures 3(a) and 3(b) show the variations of the electrical conductivity () with the temperature in the novel RTA and the novel BGK models at finite chemical potential, respectively. More precisely, we have compared our results on the electrical conductivity with those of the standard RTA and BGK models. We have also compared the result for with that for . The relaxation time appears to be momentum-dependent (eq. (16)) in the first case, whereas the relaxation time is momentum-independent in the second case. We have observed that, regardless of the relaxation type ( or ), there is a noticeable reduction in the magnitudes of in the novel RTA and the novel BGK models as compared to those in the RTA and the BGK models, respectively. This reduction in the electrical conductivity in the novel RTA and the novel BGK models is mainly attributed to the quantities and appearing in the denominators of equations (24) and (30).
T (0.16 GeV) | ||||||||
@@
In figures 4(a) and 4(b), the thermal conductivity () has been plotted as a function of the temperature in the novel RTA and the novel BGK models at finite chemical potential, respectively. These figures show the comparison of our results on the thermal conductivity with those of the standard RTA and BGK models. We have found that, irrespective of the relaxation type ( or ), there is a discernible decrease in the magnitudes of in the novel RTA and the novel BGK models as compared to those in the RTA and the BGK models, respectively. The quantities and appearing in the denominators of equations (48) and (54) mainly contribute in the decrease of the thermal conductivity in the novel RTA and the novel BGK models.
The electrical and thermal conductivities are found to be significantly impacted by the adjustments we made to the RTA model as well as the BGK model when an energy-dependent relaxation time () is used. Tables 1 and 2 show how the updated collision integral affects the charge and heat transfer in general in the low temperature and the high temperature regimes, respectively. Based on the electrical conductivity related ratios in tables 1 and 2, we can conclude that, for the novel RTA and the novel BGK models, the drop in the magnitude of becomes larger when considering the momentum dependent relaxation time as compared to the standard RTA and BGK models. We may infer from the thermal conductivity related ratios in tables 1 and 2 that, as compared to the standard RTA and BGK models, the reduction in the magnitude of for the novel RTA and the novel BGK models increases when the momentum dependent relaxation time is taken into account. Thus the momentum dependence of the relaxation time decelerates the flow of charge as well as the flow of heat in the medium.

### v3

@@ Abstract
We have studied the charge and the heat transport properties of a hot and dense QCD matter by solving the relativistic Boltzmann transport equation using a novel approximation method. Following the recently developed novel relaxation time approximation (RTA) model, we have proposed a novel Bhatnagar-Gross-Krook (BGK) model with a modified collision integral to carry out the aforementioned study. We have also compared our findings with the results of the novel RTA, the standard RTA and the standard BGK models. Our observation shows that the novel collision integrals for both the RTA and BGK models decrease the charge and the heat transport phenomena in the medium, as evidenced by the reduced values of the transport coefficients, such as the electrical conductivity and the thermal conductivity, when compared to the standard RTA and standard BGK models. Furthermore, certain observables associated with the abovementioned transport coefficients, such as the thermal diffusion constant and the Lorenz number have been explored using the novel approaches of the aforesaid models. We have found an overall decreasing trend of the thermal diffusion constant with the temperature in the novel BGK model, similar to the novel RTA model, but the magnitude remains higher throughout the temperature range. However, the magnitude of the thermal diffusion constant in the proposed novel BGK model remains conspicuously lower than its value in the standard BGK model. The magnitude of the Lorenz number in the novel BGK model remains significantly higher than that in the standard BGK model, but it is lower than that in the novel RTA model. We have also observed that the Lorenz number in all cases has an increasing trend at low temperatures, showing a violation of the Wiedemann-Franz law, whereas at high temperatures, it becomes saturated. The Lorenz number remaining above unity indicates that the thermal conductivity prevails over the electrical conductivity in the aforesaid models.
@@ 1 Introduction
A long time goal of the physics of ultrarelativistic heavy ion collisions is the observation of new states of matter predicted by the quantum chromodynamics (QCD). The experimental evidences of the relativistic heavy ion collider (RHIC) [1] at BNL and the large hadron collider (LHC) [2] at CERN indicate a new state of the strongly interacting matter known as quark-gluon plasma (QGP). Following the discovery of QGP, various properties of the strongly coupled QCD matter became a topic of high interest. Different research groups have studied different features of QGP in recent years, such as the thermodynamic properties [3, 4], the relativistic hydrodynamics [5, 6, 7], the heavy quark diffusion in QCD matter [8, 9, 10], the memory effect on heavy quark dynamics [11, 12], the dissipative spin dynamics in relativistic matter [13, 14] etc. Nowadays, there is much interest in the phase diagram region at finite temperature and baryon chemical potential [15, 16]. Many observations have also been made recently to investigate the impact of a strong magnetic field on the characteristics of hot QCD matter, such as the chiral magnetic effect [17, 18], the chiral vortical effect [19], the axial Hall current [20], the refractive indices and decay constant [21, 22], the photon and dilepton productions from QGP [23, 24, 25], the thermodynamic and magnetic properties [26, 27, 28, 29, 30, 31], the axial magnetic effect [32, 33], the thermoelectric responses of hot QCD medium in a time varying magnetic field [34, 35], the magnetic and inverse magnetic catalysis [36, 37], the nonlinear electromagnetic current [38, 39] and the magnetohydrodynamics [40, 41].
@@
It is crucial to investigate different transport coefficients since they shed light on the genesis of the hot QCD medium. Among these transport coefficients, the thermal conductivity () and the electrical conductivity () characterize the movement of heat and charges in the medium, respectively. The hydrodynamic development of the strongly interacting matter at finite baryon density also depends on these two conductivities [42, 43]. Therefore, significant effects on the observables at heavy ion collisions might result from changes in the charge and the heat transport coefficients. There are several methods for determining these conductivities, such as the relativistic Boltzmann transport equation [44, 45], the Chapman-Enskog approximation [46, 47], the lattice simulation [48, 49, 50, 51, 52, 53], the correlator technique using Green-Kubo formula [54, 55, 56] etc. Previous research has examined the impact of magnetic field on the electrical conductivity using a variety of methods, including the diluted instanton-liquid model [57, 58, 54, 59, 60, 61], the quenched SU(2) lattice gauge theory [62], the effective fugacity technique [63], the diagrammatic method utilizing the real-time formalism [64] etc. Studies on the electrical and the thermal conductivities of hot QCD matter in the presence of magnetic field within the kinetic theory approach have been carried out in references [65, 66, 67, 68].
@@
The application of theoretical methodologies intended for systems of this type has been made viable by the strongly coupled fluid behavior of QGP. A simplified version of the kinetic equation with a simplified form of the collision term is frequently used because of various technical limitations. It is referred as the relaxation time approximation (RTA) [69]. In this model, the mean-free approach reduced the complexity of the collision term. The development in the nonrelativistic case has already been achieved by Bhatnagar-Gross-Krook (BGK) model [70] and Welander model [71]. This approximation is utilized in many domains of physics and is recently used to investigate the hydrodynamization of the matter generated by the ultrarelativistic collisions of heavy ions. The relaxation time approximation has become essential for exploring the domain and microscopic basis of the relevance of hydrodynamics with relativistic effects. Although widely used, the approximation presented in ref. [69] has some fundamental challenges. It conflicts with some macroscopic and microscopic conservation laws. Recently, some works [72, 73, 74] provided a novel relaxation time approximation that fixes such fundamental issues and maintains the core features of the linearized Boltzmann collision operator. In light of the momentum-dependent relaxation time, the novel relaxation time approximation is significant as it maintains the fundamental collision features associated with the microscopic conservation laws. They showed how this adjustment impacts the particle diffusion and the bulk viscosity. In a recent study [75], the transport coefficients related to charge and heat for a rotating QGP medium have been derived within the RTA model utilizing the aforementioned collision integral.
@@
In this study, we estimate the electrical conductivity () and the thermal conductivity (), to investigate the charge and the thermal transport properties of a hot and dense QCD matter. These transport coefficients have been computed using a novel approximation framework that improves the standard RTA and standard BGK models significantly. The novel approaches of these models enhance the understanding of charge and heat transport phenomena in a strongly coupled thermal QCD medium while ensuring strict compliance with microscopic conservation laws. The findings of our study indicate that the modifications to the collision integral result in a substantial reduction of and , implying that the transport of charge and heat in QGP is more constrained than earlier assessments based on the standard RTA and standard BGK models. This can lead to revised predictions for the QGP lifetime and photon/dilepton emission rates. We also explore the thermal diffusion constant and the Lorenz number as the applications of charge and heat transport properties in an analogous environment. The thermal diffusion constant shows the overall rate of heat transport in the medium, while the Lorenz number illustrates the connection between the thermal and electrical conductivities. We use the quasiparticle description of partons, where the rest masses are replaced by the masses generated in the medium. The quark-gluon plasma is viewed as a medium made up of thermally massive quasiparticles in this framework [76, 77].
@@
The rest of the paper has been organized as follows. Using the modified collision integral, the charge and heat transport coefficients from the relativistic Boltzmann transport equation are calculated in section 2. In section 3, we have discussed the results of the aforesaid transport coefficients using the quasiparticle description of partons. Section 4 describes various applications and their results obtained from these coefficients, such as the Lorenz number and the thermal diffusion constant. In section 5, the results are summarized.
@@ 2 Charge and heat transport properties of a hot and dense QCD medium
where F is the force field acting on the particles in the medium and is the mass of the particle. The term on the right-hand side results from particle collisions within the medium. This term’s zero value denotes the Vlasov equation, which describes a collisionless system. The intricate character of the collision term presents a number of challenges even for the simplest solution of the above eq. (1). The term represents the instantaneous change in the distribution function due to collisions between particles. In order to arrive at a comprehensive solution, we will first study a few mathematical models those allow us to address the collision term in a suitable manner.
@@
where we can write and is the relaxation time for collisions, i.e. the time required to bring the perturbed system back to its equilibrium state. We can assume that the quark distribution function is near the equilibrium with a slight divergence from equilibrium in order to permit linearization of the relativistic Boltzmann transport equation. Here, the term acts as a damping frequency. The initial issue with this model is that, there is no immediate conservation of charge. To remove the crunch, the Bhatnagar-Gross-Krook (BGK) model presents a new type of collision kernel, denoted as
where represents the equilibrium particle number density and is known as the perturbed particle number density or the fluctuating particle number density. The perturbed particle number density is calculated as , which gives zero value after integrating over momentum, thus indicating instantaneous conservation of the particle number during collisions in the system. The fundamental problems regarding the nonconservation of macroscopic and microscopic conservation laws are resolved, and the essential characteristics of the linearized Boltzmann collision operator are preserved via a novel relaxation time approximation that was recently presented in a research [72].
Here, we are considering the general relativistic covariant form of the Boltzmann transport equation for an isotropic medium of quarks, antiquarks and gluons, which is given as
where is the electromagnetic field strength tensor. In order to see the effects of electric field, we take only and and vice versa components of the electromagnetic field strength tensor, i.e. and in our calculation, such that the RBTE (4) takes the following form,
@@ 2.1 Electrical conductivity in the novel RTA and the novel BGK models
The electric current density of the medium must be determined in order to see how the electric field affects the charge transport phenomena. An external electric field can cause an infinitesimal disturbance to a thermal QCD medium that contains quarks and antiquarks of distinct flavors, and gluons. Due to this electric field, there is an induced current density in the medium, which is given by
@@
Comparing eq. (6) and eq. (7), one can obtain the expression of the electrical conductivity.
By solving the relativistic Boltzmann transport equation, the infinitesimal changes in the quark and antiquark distribution functions can be obtained. The relativistic Boltzmann transport equation in the novel RTA model can be written as
@@
respectively, where , , and represents the chemical potential of flavor of quark. Getting inspired by references [78, 79, 80], we have used the relaxation time of the following form,
Here, is an arbitrary constant which controls the energy dependence of the relaxation time and has different values for different theories, for example, in QCD kinetic theories, [78] and in scalar field theories [81]. Many previously published studies on QGP transport characteristics had postulated a power-law relation between the relaxation time and the energy, with [78, 79, 80]. By including the scale dependency of the relaxation time (), we enhance the novel RTA approximation with features known in QCD. The choice of guarantees that the transport coefficients such as electrical conductivity, shear viscosity and thermal conductivity align with anticipated behaviors in quark-gluon plasma. In the analogous QCD scenario, high energy particles mainly lose energy through inelastic gluon radiation. The relaxation time scales as due to the Landau-Pomeranchuk-Migdal suppression effect. The energy loss rate of a high energy parton follows , which means that it increases with energy, but at a sublinear rate. An increased value of (e.g., ) would result in an excessively fast increase of the relaxation time with energy, hence causing an overestimation of the transport coefficients. A smaller value of (e.g., ) might make it insufficiently reliant on energy, which contradicts the anticipated quark scattering behavior in QGP. Thus, we set the value of to in the present work. In the above equation, is momentum-independent, whose expression is given [82] by
where is the QCD running coupling constant, which is a function of both temperature and chemical potential, and it has the following [83] form,
Here, is the renormalization scale parameter in the modified minimum subtraction (MS) scheme, a widely used QCD renormalization approach in the determination of the QCD running coupling constant. In case of the high temperature QGP medium, one can use GeV for one-loop running coupling constant as per lattice measurements [84] and particle data group [85]. In the above equation, the renormalization scale for electrically charged particles (quarks and antiquarks) and for gluons. Using the expression of in eq. (8), we have
@@
In the standard BGK model, the relativistic Boltzmann transport equation is written as
where the collision integral is given as
In the above equation, and represent the perturbed particle number density and the equilibrium particle number density, respectively for the flavor, which has a degeneracy of . Both can be calculated from the equations as given below,
The BGK collision kernel’s instantaneous conservation of the particle number is its main benefit over the RTA collision kernel [70, 31, 86], i.e., in case of the BGK model, we have
Furthermore, the ref. [87] had shown that, by implementing a straightforward adjustment, the BGK collision kernel effectively maintains the conservation of the color current in a covariant manner. The Boltzmann equation has been recently formulated using the BGK collision kernel to represent relativistic dissipative hydrodynamics [88].
In the present work, we proposed a collision integral in the following manner,
To fulfill the basic conservation equations in microscopic interactions, the zeroth and first moments of the collision kernel must vanish (detailed calculations are given in appendix B).
In the novel BGK model, the relativistic Boltzmann transport equation can be expressed as
Now, neglecting higher order terms , we get the solution of the above equation as
where is given by
The infinitesimal change in the quark distribution function is obtained (details are given in appendix C) as
@@ 2.2 Thermal conductivity in the novel RTA and the novel BGK models
Heat flow results from the deviation of the system from its equilibrium condition caused by the presence of a temperature gradient. Through a proportionality constant, i.e. the thermal conductivity, the heat flow vector is directly proportional to the temperature gradient. The calculation of the thermal conductivity is necessary in examining the heat transmission in a medium. In four-vector notation, the heat flow is written as
@@
where , and represent the particle number density, the energy density and the pressure, respectively, which are respectively given by the following equations,
@@
where and for very small . After solving the above equation one can extract the expression of (for a detailed calculation see appendix D) as
Similarly for antiquarks, we have
@@
In the novel BGK model, the relativistic Boltzmann transport equation takes the following form,
@@
The infinitesimal change in the quark distribution function is obtained (for a detailed calculation see appendix E) as
Similarly, for antiquark case is obtained as
@@ 3 Results and discussions
respectively, with , where the expression of is given in eq. (18). It can be seen from eq. (18) that, depends on both temperature and chemical potential. In this work, we have used GeV, corroborating consistency with the theoretical predictions and experimental data [90, 91, 92, 93]. The chosen value of the chemical potential is utilized to approximate realistic conditions in highly energetic heavy ion collisions, ensuring the numerical stability in the calculations. The accessible energy scales for a thermal QCD medium are related to the temperature, quark mass and chemical potential. The QGP phase, which is made up of light quarks, light antiquarks and gluons, has a temperature high enough to reach the maximum energy scale. The hard thermal loop (HTL) perturbation theory is one of the effective theories that is applied to tackle the divergences that arise in the computation of the QCD thermodynamic observables, the transport coefficients and the amplitudes in the high temperature regime.
In this section, the charge and the heat transport coefficients are investigated with the thermal masses as given in equations (61) and (62). In the novel RTA and the novel BGK models, quantities and appear in the denominators of the electrical conductivity and the thermal conductivity expressions. Thus the knowledge on the variations of these quantities with temperature and momentum is important to understand the variations of the conductivities. In figures 1 and 2, we have plotted and for the momentum-dependent relaxation time () and for the momentum-independent relaxation time (), respectively. We have found decreasing trends of and with the increase of temperature (figures 1(a) and 2(a)) and their increasing trends with the increase of momentum (figures 1(b) and 2(b)). In both the cases, it is observed that the momentum dependence of the relaxation time enhances both and . Thus, through the effects of the momentum-dependent relaxation time on and , it is possible to comprehend how the momentum-dependent relaxation time would modulate the transport coefficients.
Figures 3(a) and 3(b) show the variations of the electrical conductivity () with the temperature in the novel RTA and the novel BGK models at finite chemical potential, respectively. More precisely, we have compared our results on the electrical conductivity with those of the standard RTA and the standard BGK models. We have also compared the result for with that for . The relaxation time appears to be momentum-dependent (eq. (16)) in the first case, whereas the relaxation time is momentum-independent in the second case. We have observed that, regardless of the relaxation type ( or ), there is a noticeable reduction in the magnitudes of in the novel RTA and the novel BGK models as compared to those in the RTA and the BGK models, respectively. This reduction in the electrical conductivity in the novel RTA and the novel BGK models is mainly attributed to the quantities and appearing in the denominators of equations (24) and (36).
T=0.16 GeV | ||||||||
@@
In figures 4(a) and 4(b), the thermal conductivity () has been plotted as a function of the temperature in the novel RTA and the novel BGK models at finite chemical potential, respectively. These figures show the comparison of our results on the thermal conductivity with those of the standard RTA and standard BGK models. We have found that, irrespective of the relaxation type ( or ), there is a discernible decrease in the magnitudes of in the novel RTA and the novel BGK models as compared to those in the RTA and the BGK models, respectively. The quantities and appearing in the denominators of equations (54) and (60) mainly contribute in the decrease of the thermal conductivity in the novel RTA and the novel BGK models.
The electrical and thermal conductivities are found to be significantly impacted by the adjustments we made to the RTA model as well as the BGK model when an energy-dependent relaxation time () is used. Tables 1 and 2 show how the updated collision integral affects the charge and heat transfer in general in the low temperature and the high temperature regimes, respectively. Based on the electrical conductivity related ratios in tables 1 and 2, we can conclude that, for the novel RTA and the novel BGK models, the drop in the magnitude of becomes larger when considering the momentum dependent relaxation time as compared to the standard RTA and standard BGK models. We may infer from the thermal conductivity related ratios in tables 1 and 2 that, as compared to the standard RTA and standard BGK models, the reduction in the magnitude of for the novel RTA and the novel BGK models increases when the momentum dependent relaxation time is taken into account. Thus the momentum dependence of the relaxation time decelerates the flow of charge as well as the flow of heat in the medium.
On the whole, our study introduces a modified collision integral that improves the description of the electrical and the thermal transport properties in the QGP medium. Our updated collision integral considers a more accurate relaxation time dependency, which results in reduced transport coefficients, better matching the lattice QCD predictions. We have incorporated the momentum-dependent relaxation time (), capturing nontrivial scattering effects in QGP. This enables us to investigate a wider range of QCD interactions and to present a stronger suppression of charge and heat flow. Our results indicate that the QGP medium behaves as a strongly coupled fluid, exhibiting suppressed electrical and thermal conductivities in contrast to the standard RTA and standard BGK models.